# Determination of the stably free cancellation property for orders

### v1

@@ 7.8. A condition on $\mathfrak{g}$
The following statements hold.
@@ 7.9. A criterion for stable freeness
Let $\mathfrak{p}$ range over all maximal ideals of $\mathcal{O}$. For each $\mathfrak{p}$, we define an element $\alpha_{\mathfrak{p}}\in X_{\mathfrak{p}}$ with $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}$ as follows. If $\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ then we set $\alpha_{\mathfrak{p}}:=\beta$. If $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$, we let $\alpha_{\mathfrak{p}}^{\prime}\in X_{\mathfrak{p}}$ be an arbitrarily chosen element such that $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}^{\prime}$. Then $\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}^{\prime}+\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}$, so $\overline{\alpha}_{\mathfrak{p}}^{\prime}\in\left(\Lambda_{\mathfrak{p}}/\mathfrak{f}_{\mathfrak{p}}\right)^{\times}$. Since $\Lambda_{\mathfrak{p}}$ is semilocal, the canonical map $\Lambda_{\mathfrak{p}}^{\times}\longrightarrow(\Lambda_{\mathfrak{p}}/\mathfrak{f}_{\mathfrak{p}})^{\times}$ is surjective by [Bas68, Chapter III, (2.9) Corollary]. In particular, there exists $\alpha_{\mathfrak{p}}^{\prime\prime}\in\Lambda_{\mathfrak{p}}^{\times}$ such that $\alpha_{\mathfrak{p}}^{\prime}\equiv\alpha_{\mathfrak{p}}^{\prime\prime}\pmod{\mathfrak{f}_{\mathfrak{p}}}$. Then we set $\alpha_{\mathfrak{p}}:=(\alpha_{\mathfrak{p}}^{\prime\prime})^{-1}\alpha_{\mathfrak{p}}^{\prime}$ and note that $\alpha_{\mathfrak{p}}\equiv 1\pmod{\mathfrak{f}_{\mathfrak{p}}}$.
@@
Then the class of
$\mathfrak{a}=\prod_{\mathfrak{P}}\mathfrak{P}^{v_{\mathfrak{P}}(\gamma_{\mathfrak{P}}\xi)}$
is a representative of $\psi(\gamma\bmod L^{\times+}\mathrm{nr}(U_{\mathfrak{f}}(\mathcal{M})))$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$. In the definition of $\mathfrak{a}$ the product extends over all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{L}$.
Using the canonical isomorphism of (10.2) we write $\alpha_{\mathfrak{p}}=\left(\alpha_{\mathfrak{P}}\right)_{\mathfrak{P}\mid\mathfrak{p}}$ with $\alpha_{\mathfrak{P}}\in A_{\mathfrak{P}}$ for the local basis elements $\alpha_{\mathfrak{p}}$ chosen as above. Then, for $\mathfrak{P}\mid\mathfrak{g}$, the result of [BB06, Corollary 2.4] implies $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})\equiv 1\pmod{\mathfrak{g}_{\mathfrak{P}}}$, so that we can and do choose $\xi=1$.
@@
Hence there exists an element $\gamma\in\mathcal{M}_{\mathfrak{p}}$ such that $\gamma\beta\equiv 1\pmod{\mathfrak{f}_{\mathfrak{p}}}$. It follows that $\gamma\beta\equiv 1\pmod{\mathfrak{f}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. From [BB06, Corollary 2.4] we obtain $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\gamma\beta)\equiv 1\pmod{\mathfrak{g}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. Since $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$ by Lemma 7.9, assumption (G) implies that $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$. Hence $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))=0$. Since $\alpha_{\mathfrak{p}}\equiv 1\pmod{\mathfrak{f}_{\mathfrak{p}}}$ the same argument shows that $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))=0$ for all maximal ideals $\mathfrak{P}\mid\mathfrak{p}$.
@@ 7.10. Choosing $\mathfrak{g}$ first
Recall that in §7.1 and §7.4, respectively, we first chose $\mathfrak{f}$ to be a proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and then set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$. In some applications, it is convenient to work the other way round: first choose an proper full ideal $\mathfrak{g}$ of $\mathcal{O}_{C}$ such that $\mathfrak{g}\mathcal{M}\subseteq\Lambda$ and then set $\mathfrak{f}:=\mathfrak{g}\mathcal{M}$. The following lemma shows that the latter approach is consistent with the former.
@@
Let $\mathfrak{g}$ be a proper full ideal of $\mathcal{O}_{C}$ such that $\mathfrak{f}:=\mathfrak{g}\mathcal{M}\subseteq\Lambda$. Then $\mathfrak{f}$ is a proper full two sided-ideal of $\mathcal{M}$ and $\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{g}$.
@@
The first claim is clear. To prove the second claim, we can and do work componentwise and hence assume without loss of generality that $A=\mathrm{Mat}_{n}(D)$ for some $n\geq 1$ and some skewfield $D$ containing $K$. Let $F=Z(D)$ and note that $F=C$.
Let $\Delta$ be any maximal $\mathcal{O}_{F}$-order in $D$. Then $\Delta\cap F=\mathcal{O}_{F}$. Thus if $\alpha\delta\in\mathcal{O}_{F}$ for some $\delta\in\Delta$ and some $\alpha\in\mathcal{O}_{F}$ with $\alpha\neq 0$ then $\delta=\alpha^{-1}(\alpha\delta)\in\Delta\cap F=\mathcal{O}_{F}$. This implies that as an $\mathcal{O}_{F}$-module, $\Delta/\mathcal{O}_{F}$ is torsion-free and hence is projective. Thus the canonical short exact sequence of $\mathcal{O}_{F}$-modules
@@
$\mathfrak{g}\Delta\cap\mathcal{O}_{F}=\mathfrak{g}\mathcal{O}_{F}\oplus\mathfrak{g}(\Delta/\mathcal{O}_{F})\cap\mathcal{O}_{F}=\mathfrak{g}.$ (7.12)
@@ 8.1. Generalised Swan modules
For $\beta\in\mathcal{M}\cap A^{\times}$ we define $X_{\beta}=X_{{\beta},\Lambda,\mathcal{M}}=\mathcal{M}\beta\cap\Lambda$, and call this the generalised Swan module attached to $\beta$ (and $\Lambda,\mathcal{M}$). Note that $X_{\beta}$ is a left ideal of $\Lambda$.
The following lemma determines the localisations of certain generalised Swan modules.
@@
Now suppose that $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. Then $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ and so by Assumption (G) and Lemma 7.9 and we have that $\mathfrak{f}_{\mathfrak{P}}\subseteq\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$. Then $\mathcal{M}_{\mathfrak{P}}\beta+\mathfrak{f}_{\mathfrak{P}}=(\mathcal{M}+\mathfrak{f})_{\mathfrak{P}}=\mathcal{M}_{\mathfrak{P}}$ and so Nakayama’s lemma implies that $\mathcal{M}_{\mathfrak{P}}\beta=\mathcal{M}_{\mathfrak{P}}$ for all $\mathfrak{P}\mid\mathfrak{p}$, or equivalently, $\beta\in\mathcal{M}_{\mathfrak{p}}^{\times}$. Hence $\mathcal{M}_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}$ and so
@@
The following result establishes further properties of certain generalised Swan modules. We recall that the maps $\pi,\mu$ and $\tilde{\nu}$ were defined in (7.1), (7.8) and (7.9), respectively.
@@
    $X_{\beta}$ is free over $\Lambda$ if and only if there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{\beta})=\pi(\overline{u})$.
@@
Asserion (v) is just the definition of $\tilde{\nu}$ (7.9). Assertion (vi) follows from (ii), (iii) and (iv) together with Theorem 7.10. We have that $X_{\beta}$ is stably free over $\Lambda$ if and only if $[X_{\beta}]_{\mathrm{st}}=0$ and by (7.10) we have $\ker(\mu)=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$. Therefore assertion (vii) now follows from (vi). Finally, assertion (viii) follows from (iii) and (iv) together with Corollary 7.3. ∎
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta_{1},\beta_{2}\in\mathcal{M}\cap A^{\times}$ such that $\overline{\beta_{1}},\overline{\beta_{2}}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Suppose that $\overline{\beta_{1}},\overline{\beta_{2}}$ have the same image in ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Then
@@
    $X_{\beta_{1}}$ is free over $\Lambda$ if and only if $X_{\beta_{2}}$ is free over $\Lambda$, and
@@ 8.2. An algorithm for determining SFC
    Compute $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ and $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$ as finite abelian groups.
@@
In step (3), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. If step (4) is reached then $\Lambda$ must be non-maximal. Note that $\mathfrak{f}$ as computed in step (4) is a two-sided ideal of $\mathcal{M}$ such that $\mathfrak{f}\mathcal{M}\subseteq\Lambda$, and so $\mathfrak{f}$ must be proper. In step (11), the correctness of the output follows from Theorem 8.7. ∎
@@
Steps (9) and (10): Since both rings $\mathcal{M}/\mathfrak{f}$ and $\Lambda/\mathfrak{f}$ are finite, their units can be computed as sets in finite time. (For a more practical approach, see §11). Given $\beta\in S$, algorithms for finitely generated abelian groups, as presented for example in [Coh00, §4.1], can be used to determine whether $\tilde{\nu}(\overline{\beta})\in\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$.
Step (11): Given $\beta\in T$, [BJ11, Algorithm 3.1] can be used to determine whether $X_{\beta}$ is free over $\Lambda$ (though stated for orders in group algebras, this algorithm can also be applied to orders in arbitrary finite-dimensional semisimple algebras as explained in [BJ11, Remark 4.9]). Note that this algorithm requires certain hypotheses, which ensure that one can find the preimage of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$, as well as a generator of $\mathcal{M}X_{\beta}$ as an $\mathcal{M}$-module. Since generators of $\mathcal{M}^{\times}$ can in principle be found using the algorithms of [BCNS15], and since $\mathcal{M}X_{\beta}=\mathcal{M}\beta$ by Proposition 8.3 (iii), these hypotheses are in fact unnecessary in this situation. ∎
@@
Then by Lemma 7.11 the ideals $\mathfrak{f}$ and $\mathfrak{g}$ would satisfy $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{f}\cap C$. However, the disadvantage of this approach is that in general it will produce a smaller choice of $\mathfrak{f}$ and thus a larger choices of the sets $S$ and $T$.

### v2

@@ 7.8. A condition on $\mathfrak{g}$
The following statements hold:
@@ 7.9. A criterion for stable freeness
Let $\mathfrak{p}$ range over all maximal ideals of $\mathcal{O}$. For each $\mathfrak{p}$, we define an element $\alpha_{\mathfrak{p}}\in X_{\mathfrak{p}}$ with $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}$ as follows. If $\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}=\mathcal{M}_{\mathfrak{p}}$ then we set $\alpha_{\mathfrak{p}}:=\beta$. If $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$, we let $\alpha_{\mathfrak{p}}^{\prime}\in X_{\mathfrak{p}}$ be an arbitrarily chosen element such that $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}^{\prime}$. Then $\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}^{\prime}+\mathfrak{f}_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}$, so $\overline{\alpha}_{\mathfrak{p}}^{\prime}\in\left(\Lambda_{\mathfrak{p}}/\mathfrak{f}_{\mathfrak{p}}\right)^{\times}$. Since $\Lambda_{\mathfrak{p}}$ is semilocal, the canonical map $\Lambda_{\mathfrak{p}}^{\times}\longrightarrow(\Lambda_{\mathfrak{p}}/\mathfrak{f}_{\mathfrak{p}})^{\times}$ is surjective by [Bas68, Chapter III, (2.9) Corollary]. In particular, there exists $\alpha_{\mathfrak{p}}^{\prime\prime}\in\Lambda_{\mathfrak{p}}^{\times}$ such that $\alpha_{\mathfrak{p}}^{\prime\prime}\equiv\alpha_{\mathfrak{p}}^{\prime}\bmod{\mathfrak{f}_{\mathfrak{p}}}$. Then we set $\alpha_{\mathfrak{p}}:=(\alpha_{\mathfrak{p}}^{\prime\prime})^{-1}\alpha_{\mathfrak{p}}^{\prime}$ and note that $\alpha_{\mathfrak{p}}\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$.
@@
If $\mathfrak{a}=\prod_{\mathfrak{P}}\mathfrak{P}^{v_{\mathfrak{P}}(\gamma_{\mathfrak{P}}\xi)},$ where the product extends over all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{L}$, then the class of $\mathfrak{a}$ is a representative of $\psi(\gamma\bmod L^{\times+}\mathrm{nr}(U_{\mathfrak{f}}(\mathcal{M})))$ in $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$.
Using the canonical isomorphism of (7.7) we write $\alpha_{\mathfrak{p}}=\left(\alpha_{\mathfrak{P}}\right)_{\mathfrak{P}\mid\mathfrak{p}}$ with $\alpha_{\mathfrak{P}}\in A_{\mathfrak{P}}$ for the local basis elements $\alpha_{\mathfrak{p}}$ chosen as above. We will apply the above recipe for $\gamma_{\mathfrak{P}}=\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})$. The result of [BB06, Corollary 2.4] implies that for $\mathfrak{P}\mid\mathfrak{g}$, we have $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})\equiv 1\bmod{\mathfrak{g}_{\mathfrak{P}}}$, so we can and do choose $\xi=1$.
@@
Hence there exists an element $\gamma\in\mathcal{M}_{\mathfrak{p}}$ such that $\gamma\beta\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$. It follows that $\gamma\beta\equiv 1\bmod{\mathfrak{f}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. From [BB06, Corollary 2.4] we obtain $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\gamma\beta)\equiv 1\bmod{\mathfrak{g}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. Since $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$ by Lemma 7.9, assumption (G) implies that $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$. Hence $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))=0$. Since $\alpha_{\mathfrak{p}}\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$ the same argument shows that $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))=0$ for all maximal ideals $\mathfrak{P}\mid\mathfrak{p}$.
@@ 7.10. Choosing $\mathfrak{g}$ first
Recall that in §7.1 and §7.4, respectively, we first chose $\mathfrak{f}$ to be a proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and then set $\mathfrak{g}:=\mathfrak{f}\cap\mathcal{O}_{C}$. In some applications, it is convenient to work the other way round: first choose a proper full ideal $\mathfrak{g}$ of $\mathcal{O}_{C}$ such that $\mathfrak{g}\mathcal{M}\subseteq\Lambda$ and then set $\mathfrak{f}:=\mathfrak{g}\mathcal{M}$. The following lemma shows that the latter approach is consistent with the former.
@@
Let $\mathfrak{g}$ be a proper full ideal of $\mathcal{O}_{C}$ such that $\mathfrak{f}:=\mathfrak{g}\mathcal{M}\subseteq\Lambda$. Then $\mathfrak{f}$ is a proper full two-sided ideal of $\mathcal{M}$ and $\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{g}$.
@@
The first claim is clear. To prove the second claim, we can and do work componentwise and hence assume without loss of generality that $A=\mathrm{Mat}_{n}(D)$ for some $n\geq 1$ and some division algebra $D$ containing $K$. Let $F=Z(D)$ and note that $F=C$.
Let $\Delta$ be any maximal $\mathcal{O}_{F}$-order in $D$. Then $\Delta\cap F=\mathcal{O}_{F}$. Thus if $\alpha\delta\in\mathcal{O}_{F}$ for some $\delta\in\Delta$ and some $\alpha\in\mathcal{O}_{F}$ with $\alpha\neq 0$ then $\delta=\alpha^{-1}(\alpha\delta)\in\Delta\cap F=\mathcal{O}_{F}$. This implies that as an $\mathcal{O}_{F}$-module, $\Delta/\mathcal{O}_{F}$ is torsion-free and hence projective. Thus the canonical short exact sequence of $\mathcal{O}_{F}$-modules
@@
$\mathfrak{g}\Delta\cap\mathcal{O}_{F}=\left(\mathfrak{g}\mathcal{O}_{F}\oplus\mathfrak{g}(\Delta/\mathcal{O}_{F})\right)\cap\mathcal{O}_{F}=\mathfrak{g}.$ (7.12)
@@ 8.1. Test lattices
For $\beta\in\mathcal{M}\cap A^{\times}$ we define $X_{\beta}=X_{{\beta},\Lambda,\mathcal{M}}=\mathcal{M}\beta\cap\Lambda$, and call this the test lattice attached to $\beta$ (and $\Lambda,\mathcal{M}$). Note that $X_{\beta}$ is a left ideal of $\Lambda$.
The following lemma determines the localisations of certain test lattices.
@@
Now suppose that $\Lambda_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$. Then $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}$ and so by Assumption (G) and Lemma 7.9 we have that $\mathfrak{f}_{\mathfrak{P}}\subseteq\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$. Thus
$\mathcal{M}_{\mathfrak{P}}\beta+\mathfrak{f}_{\mathfrak{P}}=(\mathcal{M}+\mathfrak{f})_{\mathfrak{P}}=\mathcal{M}_{\mathfrak{P}}$
and so Nakayama’s lemma implies that $\mathcal{M}_{\mathfrak{P}}\beta=\mathcal{M}_{\mathfrak{P}}$ for all $\mathfrak{P}\mid\mathfrak{p}$, or equivalently, $\beta\in\mathcal{M}_{\mathfrak{p}}^{\times}$. Hence $\mathcal{M}_{\mathfrak{p}}\beta=\mathcal{M}_{\mathfrak{p}}$ and so
@@
The following result establishes further properties of certain test lattices. We recall that the maps $\pi,\mu$ and $\tilde{\nu}$ were defined in (7.1), (7.8) and (7.9), respectively.
@@
    $X_{\beta}$ is free over $\Lambda$ if and only if there exists $u\in\mathcal{M}^{\times}$ such that $\pi(\overline{u})=\pi(\overline{\beta})$.
@@
Assertion (v) is just the definition of $\tilde{\nu}$ (7.9). Assertion (vi) follows from (ii), (iii) and (iv) together with Theorem 7.10. We have that $X_{\beta}$ is stably free over $\Lambda$ if and only if $[X_{\beta}]_{\mathrm{st}}=0$ and by (7.10) we have $\ker(\mu)=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$. Therefore assertion (vii) now follows from (vi). Finally, assertion (viii) follows from (iii) and (iv) together with Corollary 7.3. ∎
@@
Suppose that $\mathfrak{g}$ satisfies Assumption (G). Let $\beta_{1},\beta_{2}\in\mathcal{M}\cap A^{\times}$ such that $\overline{\beta_{1}},\overline{\beta_{2}}\in(\mathcal{M}/\mathfrak{f})^{\times}$. Suppose that $\overline{\beta_{1}},\overline{\beta_{2}}$ have the same image in ×\$\mathcal{M}^{\times}\backslash(\mathcal{M}/\mathfrak{f})^{\times}{}{/}{}(\Lambda/\mathfrak{f})^{\times}$. Then the following statements hold:
@@
    $X_{\beta_{1}}$ is free over $\Lambda$ if and only if $X_{\beta_{2}}$ is free over $\Lambda$.
@@ 8.2. An algorithm for determining SFC
    Compute $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ and $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})\leq I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ as finite abelian groups.
@@
In step (3), the correctness of an output $\mathsf{false}$ follows from Corollary 3.2 and the correctness of an output $\mathsf{true}$ is clear. If step (4) is reached then $\Lambda$ must be non-maximal. Note that $\mathfrak{f}$ as computed in step (4) is a two-sided ideal of $\mathcal{M}$ such that $\mathfrak{f}\mathcal{M}\subseteq\Lambda$, and so $\mathfrak{f}$ must be proper. By Proposition 8.3 (v),(vii) we have $T=\{\beta\in S\colon X_{\beta}\text{ is stably free over }\Lambda\}$. Thus in step (11), the correctness of the output follows from Theorem 8.7. ∎
@@
Steps (9) and (10): Since both rings $\mathcal{M}/\mathfrak{f}$ and $\Lambda/\mathfrak{f}$ are finite, their units can be computed as sets in finite time. (For a more practical approach, see §11). Given $\beta\in S$, algorithms for finitely generated abelian groups, as presented for example in [Coh00, §4.1], can be used to determine whether $\tilde{\nu}(\overline{\beta})=(\mathrm{nr}(\beta)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})$ belongs to $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$.
Step (11): Given $\beta\in T$, [BJ11, Algorithm 3.1] (see also [HJ20, Algorithm 4.1]) can be used to determine whether $X_{\beta}$ is free over $\Lambda$ (though stated for orders in group algebras, this algorithm can also be applied to orders in arbitrary finite-dimensional semisimple algebras as explained in [BJ11, Remark 4.9]). Note that this algorithm requires certain hypotheses, which ensure that one can find the preimage of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$, as well as a generator of $\mathcal{M}X_{\beta}$ as an $\mathcal{M}$-module. Since generators of $\mathcal{M}^{\times}$ can in principle be found using the algorithms of [BCNS15], and since $\mathcal{M}X_{\beta}=\mathcal{M}\beta$ by Proposition 8.3 (iii), these hypotheses are in fact unnecessary in this situation. Note that we also give an alternative freeness testing algorithm in §12 that may be more efficient under certain circumstances. ∎
@@
Then by Lemma 7.11 the ideals $\mathfrak{f}$ and $\mathfrak{g}$ would satisfy $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}=\mathfrak{f}\cap C$. However, the disadvantage of this approach is that in general it will produce a smaller choice of $\mathfrak{f}$ and thus larger choices of the sets $S$ and $T$.